# Kővári-Sós-Turán theorem for hereditary families

### v1

@@ 2.3 Finishing the proof
then $G$ contains a subset $X$ of at least $\ell$ vertices such that every $k$ vertices in $U$ have at least $m$ common neighbors.
@@ 2.4 Lower bounds
Let $H=K_{k,\ell}$ where $\ell\geq(k-1)!$. Then, for all integers $n\geq s\geq 2k!$ we have
@@
    Finally, one can argue similarly that if $G$ contains $K_{s,s}$ as a subgraph, then $G_{0}$ must contain $K_{k!,k!}$, contradiction. This completes the proof. ∎
One can slightly improve the above construction when $n=2^{O(s)}$ by replacing vertices in the blow-up by Ramsey graphs avoiding $K_{s,s}$ and independent sets of size $h$, instead of replacing them by cliques. In this regime, the construction can be improved to show that $\mathrm{ex}^{*}(n,H,s)\geq s^{|V(H)|^{1/(k+1)}}n^{2-\frac{1}{k}}$, i.e., that the exponent of $s$ should grow with $|V(H)|$. We omit further details since there is still a large gap between upper and lower bounds.
@@ 3 Trees, induced $C_{4}$-free subgraphs and the Erdős-Hajnal conjecture
We note that the sufficiently large condition in the above result is not especially quantitative. It would be interesting to determine whether we can take $d\leq k^{O(1/\varepsilon)}$.
@@
    Apply Proposition 3.1 to the graph $G$ with parameters $H:=T$, $k:=4t+1$ and $\varepsilon:=\frac{1}{4t}$. We conclude that, for all sufficiently large $s$, either $G$ contains an induced $C_{4}$-free subgraph $G^{\prime}\subseteq G$ with average degree at least $4t$, or an induced subgraph $G^{\prime}\subseteq G$ on $n^{\prime}\geq s^{Ct/10}$ vertices and average degree $(n^{\prime})^{1-\frac{1}{4t}}$.
@@
    By our assumption on $G$, the first conclusion of the Proposition 3.1 cannot hold and therefore $G$ contains an induced subgraph $G^{\prime}\subseteq G$ on $n^{\prime}\geq d(G)^{1/10}\geq s^{100k^{2}}>(C_{H_{k}}s)^{8|V(H_{k})|+20}$ vertices and with $e(G^{\prime})\geq(n^{\prime})^{1-\frac{1}{8k}}$ edges. Since $H_{k}$ is a $(q+1)$-regular graph, its degrees are bounded by $2k$ and therefore part (ii) of Theorem 2.1 shows that $G^{\prime}$ contains an induced copy of $H_{k}$. This presents a contradiction to the assumption that $G$ has no $C_{4}$-free induced subgraph of average degree at least $k$, thus completing the proof. ∎
@@ 4 Cycles and the cube
The goal of this section is to prove Theorem 1.4, which is concerned with graphs that are induced $C_{2k}$-free and $K_{s,s}$-free. Also, we show Proposition 1.5 which gives an upper bound on the induced Turán number of the cube graph. We begin by proving the following strengthening of Theorem 1.4, which is needed to for our proof of Theorem 1.5.
@@
Let $\alpha>0$ be a fixed real number and let $K=2^{3/\alpha+4}$. Furthermore, let $G$ be a graph on $n$ vertices with at least $Cn^{1+\alpha}$ edges. Then, there exists a $K$-almost-regular induced subgraph $H\subseteq G$ on $m$ vertices with at least $\frac{C}{4}m^{1+\alpha}$ edges.
@@
    where the last inequality follows form the choice of $K_{0}$.
    Thus, we conclude that $U$ is incident to at most $e(G)/2$ edges and so $V(G)\backslash U$ induces at least $e(G)/2$ edges. Let us now focus on the subgraph $G^{\prime}\subseteq G$ induced on the set $V(G)\backslash U$. We may remove from this graph all vertices of degree at most $d(G^{\prime})/2$, while removing at most half of the edges of $G^{\prime}$. In this way, we obtain an induced subgraph $G^{\prime\prime}\subseteq G$ with the property that $e(G^{\prime\prime})\geq e(G^{\prime})/2\geq e(G)/4$ and $\delta(G^{\prime\prime})\geq d(G^{\prime})/2\geq d(G)/4$. Since all vertices of $G^{\prime}$ have degree at most $K_{0}d(G)$, we finally arrive at the conclusion $4K_{0}\delta(G^{\prime\prime})\geq\Delta(G^{\prime\prime})$, thus completing the proof. ∎
@@
Let $k$ be a positive integer, let $K=2^{3k+4}$ and let $G$ be a $K$-almost-regular graph with $d(G)=Cn^{1/k}$. Then at most $\frac{2^{2k+10}}{\sqrt{C}}\hom(C_{2k},G)$ homomorphic $2k$-cycles in $G$ are degenerate.
@@
    $=\frac{c^{3}\delta^{-2}}{200}+\frac{c^{2}\delta^{-1}}{20}+\frac{c^{3}\delta^{-2}}{1000}+\frac{2c^{2}\delta^{-1}}{100}+t^{2}<\frac{c^{3}\delta^{-2}}{100}=e(G_{B})p^{2},$
@@
    Let us begin by applying Lemma 4.2 with $\alpha=\frac{1}{k}$ to the graph $G_{1}$ of edges crossing the partition $(A,B)$. Let $C_{0}=\frac{C}{4}$ and $K=2^{3k+4}$, then there is a $K$-almost-regular induced subgraph $H_{1}$ of $G_{1}$ satisfying $e(H_{1})\geq C_{0}m^{1+1/k}$, where $m$ is the number of vertices of $H_{1}$. Also, let $H$ be the subgraph of $G$ induced by $V(H_{1})$.
    For any two distinct vertices $u,v\in V(H)$, we denote by $\mathcal{P}_{u,v}$ the set of alternating paths of length $k$ between vertices $u$ and $v$ in $H$, and let $P_{u,v}=|\mathcal{P}_{u,v}|$. Furthermore, let $A_{u,v}$ denote the number of ordered pairs of paths $(P_{1},P_{2})\in\mathcal{P}_{u,v}^{2}$ which intersect in a vertex different from $u,v$. Finally, let $B_{u,v}$ denote the number of ordered pairs of paths $(P_{1},P_{2})\in\mathcal{P}_{u,v}^{2}$ such that some internal vertex of $P_{1}$ and some internal vertex of $P_{2}$ are connected by an edge of $H$.
@@
    Finally, if $H$ has no induced nondegenerate $2k$-cycles, then $\sum_{u,v}B_{u,v}\geq\frac{1}{4k}\hom(C_{2k},H_{1})$. To see why, note that every nondegenerate cycle $(x_{1},\dots,x_{2k})$ must have a chord, since it is not induced. Then, there exists a way to shift the indices of the vertices, say by setting $y_{i}=x_{i+t\bmod 2k}$ for some $t$, such that the chord goes between the vertex sets $\{y_{2},\dots,y_{k-1}\}$ and $\{y_{k+2},\dots,y_{2k}\}$. In this case, the cycle $(y_{1},\dots,y_{2k})$ corresponds to a pair of paths between $y_{1}$ and $y_{k+1}$ with an edge between their internal vertices, which is counted in $B_{y_{1},y_{k+1}}$. This argument shows that for any non-degenerate $2k$-cycle, one of its $2k$ cyclic relabeling corresponds to a pair of paths counted in the sum $\sum_{u,v}B_{u,v}$. Hence, $2k\sum_{u,v}B_{u,v}$ is at least the number of nondegenerate $2k$-cycles, and so $\sum_{u,v}B_{u,v}\geq\frac{1}{4k}\cdot\hom(C_{2k},H_{1})$.
@@
    Let us fix these vertices and define $\delta=2^{2k+10}\cdot{8k}/{\sqrt{C_{0}}}$. Since $B_{u,v}$ counts the number of pairs of paths in $\mathcal{P}_{u,v}$, we always have $B_{u,v}\leq P_{u,v}^{2}$. Thus, we deduce that $8k\delta^{-1}A_{u,v}\leq 8kB_{u,v}\leq 8kP_{u,v}^{2}$ and so $A_{u,v}\leq\delta P_{u,v}^{2}$. Also, we directly have both $B_{u,v}\geq\frac{1}{16k}P_{u,v}^{2}$ and $8kP_{u,v}^{2}\geq 8kB_{u,v}\geq\frac{1}{2}C_{0}^{2k}$, implying $P_{u,v}\geq C_{0}^{k}/4\sqrt{k}$.
    We now define two auxiliary graphs, a red and a blue one, on the set of vertices $\mathcal{P}_{u,v}$. A pair of distinct paths $P_{1},P_{2}\in\mathcal{P}_{u,v}$ is a red edge if the paths $P_{1},P_{2}$ intersect in some of their internal vertices. Also, $P_{1}P_{2}$ is a blue edge if there exists an edge of $H$ between the internal vertices of $P_{1}$ and $P_{2}$. Let us denote the red graph by $G_{R}$ and the blue graph by $G_{B}$. Note that an ordered pair of paths $(P_{1},P_{2})$ is counted in $A_{u,v}$ precisely if $P_{1}P_{2}$ is a red edge or if $P_{1}=P_{2}$. Thus, we have that the number of red edges is precisely $e(G_{R})=(A_{u,v}-P_{u,v})/{2}\leq\delta P_{u,v}^{2}$. Similarly, we have $e(G_{B})=(B_{u,v}-P_{u,v})/{2}\geq\frac{1}{32k}P_{u,v}^{2}$.
@@ 4.1 The cube graph
We finish the section by showing Proposition 1.5, which states that $\mathrm{ex}^{*}(n,Q_{8},s)\leq O_{s}(n^{8/5})$, where $Q_{8}$ denotes for the graph of a three-dimensional cube. Before we start the proof, we show that almost all paths of a given length in a $K_{s,s}$-free graph are induced. Although we need this statement only for paths of length three, we prove it in full generality since we believe it might be of independent interest.
@@
    Let $G$ be an $n$ vertex graph with average degree at least $Cn^{3/5}$ containing no $K_{s,s}$. Partition the vertex set of $G$ into parts $A,B$ such that at least half of the edges cross the partition, and let $G_{1}$ be the graph of crossing edges. By Lemma 4.2 applied with $\alpha=\frac{3}{5}$, $G_{1}$ has a $K$-almost-regular induced subgraph $H_{1}$ with $m$ vertices and average degree $d$ satisfying $d\geq C_{0}m^{3/5}$, where $C_{0}=\frac{C}{4}$ and $K=2^{9}$. Let $H$ be the subgraph of $G$ induced on $V(H_{1})$.
@@
    We claim that if $C_{0}$ is sufficiently large, one has $e(U,V)\geq C_{0}(|U|+|V|)^{4/3}$, which allows us to apply Proposition 4.1 with $k=3$. To verify that there are sufficiently many edges between $U$ and $V$, we have
    $e(U,V)=\Omega(d^{3}/n)\geq\Omega(C_{0}^{3}n^{4/5})\geq C_{0}\cdot(2Kd)^{4/3}\geq C_{0}(|U|+|V|)^{4/3}.$

### v2

@@ 2.3 Finishing the proof
then $G$ contains a subset $X$ of at least $\ell$ vertices such that every $k$ vertices in $X$ have at least $m$ common neighbors.
@@ 2.4 Lower bounds
Let $H=K_{k,\ell}$ where $\ell=(k-1)!$. Then, for all integers $n\geq s\geq 2k!$ we have
@@
    Finally, one can argue that if $G$ contains $K_{s,s}$ as a subgraph, then $G_{0}$ must contain $K_{s/t,s/t}=K_{2k!,2k!}$, contradiction. This completes the proof. ∎
One can slightly improve the above construction when $n=2^{O(s)}$ by replacing vertices in the blow-up by Ramsey graphs avoiding $K_{s/2,s/2}$ and independent sets of size $h$, instead of replacing them by cliques. If one replaces the edges between these vertices with dense graphs which avoid $K_{s/2,s/2}$, the resulting graph has no $K_{s,s}$. Hence, this improved construction shows that $\mathrm{ex}^{*}(n,H,s)\geq s^{|V(H)|^{1/(k+1)}}n^{2-\frac{1}{k}}$, i.e., that the exponent of $s$ should grow with $|V(H)|$. We omit further details since there is still a large gap between upper and lower bounds.
@@ 3 Trees, induced $C_{4}$-free subgraphs and the Erdős-Hajnal conjecture
We note that the sufficiently large condition in the above result is not especially quantitative. It would be interesting to determine whether we can take $d\leq k^{O_{H}(1/\varepsilon)}$.
@@
    Apply Proposition 3.1 to the graph $G$ with parameters $H:=T$, $k:=4t+1$ and $\varepsilon:=\frac{1}{4t}$. We conclude that, for all sufficiently large $s$, either $G$ contains an induced $C_{4}$-free subgraph $G^{\prime}\subseteq G$ with average degree greater than $4t$, or an induced subgraph $G^{\prime}\subseteq G$ on $n^{\prime}\geq s^{Ct/10}$ vertices and average degree $(n^{\prime})^{1-\frac{1}{4t}}$.
@@
    By our assumption on $G$, the first conclusion of the Proposition 3.1 cannot hold and therefore $G$ contains an induced subgraph $G^{\prime}\subseteq G$ on $n^{\prime}\geq d(G)^{1/10}\geq s^{100k^{2}}>(C_{H_{k}}s)^{8|V(H_{k})|+20}$ vertices and with $e(G^{\prime})\geq(n^{\prime})^{2-\frac{1}{8k}}$ edges. Since $H_{k}$ is a $(q+1)$-regular graph, its degrees are bounded by $2k$ and therefore part (ii) of Theorem 2.1 shows that $G^{\prime}$ contains an induced copy of $H_{k}$. This presents a contradiction to the assumption that $G$ has no $C_{4}$-free induced subgraph of average degree at least $k$, thus completing the proof. ∎
@@ 4 Cycles and the cube
The goal of this section is to prove Theorem 1.4, which is concerned with graphs that are induced $C_{2k}$-free and $K_{s,s}$-free. Also, we show Proposition 1.5 which gives an upper bound on the induced Turán number of the cube graph. We begin by proving the following strengthening of Theorem 1.4, which is needed to for our proof of Proposition 1.5.
@@
Let $\alpha>0$ be a fixed real number and let $K=2^{3/\alpha+5}$. Furthermore, let $G$ be a graph on $n$ vertices with at least $Cn^{1+\alpha}$ edges. Then, there exists a $K$-almost-regular induced subgraph $H\subseteq G$ on $m$ vertices with at least $\frac{C}{4}m^{1+\alpha}$ edges.
@@
    where the last inequality follows from the choice of $K_{0}$.
    Thus, we conclude that $U$ is incident to at most $e(G)/2$ edges and so $V(G)\backslash U$ induces at least $e(G)/2$ edges. Let us now focus on the subgraph $G^{\prime}\subseteq G$ induced on the set $V(G)\backslash U$. We may repeatedly remove from this graph all vertices of degree at most $d(G^{\prime})/4$, while removing at most half of the edges of $G^{\prime}$. In this way, we obtain an induced subgraph $G^{\prime\prime}\subseteq G$ with the property that $e(G^{\prime\prime})\geq e(G^{\prime})/2\geq e(G)/4$ and $\delta(G^{\prime\prime})\geq d(G^{\prime})/4\geq d(G)/8$. Since all vertices of $G^{\prime}$ have degree at most $K_{0}d(G)$, we finally arrive at the conclusion $8K_{0}\delta(G^{\prime\prime})\geq\Delta(G^{\prime\prime})$, thus completing the proof. ∎
@@
Let $k$ be a positive integer, let $K=2^{3k+5}$ and let $G$ be a $K$-almost-regular graph with $d(G)=Cn^{1/k}$. Then at most $\frac{2^{2k+10}}{\sqrt{C}}\hom(C_{2k},G)$ homomorphic $2k$-cycles in $G$ are degenerate.
@@
    $=\frac{c^{3}\delta^{-2}}{200}+\frac{c^{2}\delta^{-1}}{20}+\frac{c^{3}\delta^{-2}}{1000}+\frac{2c^{2}\delta^{-1}}{100}+t^{2}<\frac{c^{3}\delta^{-2}}{100}\leq e(G_{B})p^{2},$
@@
    Let us begin by applying Lemma 4.2 with $\alpha=\frac{1}{k}$ to the graph $G_{1}$ of edges crossing the partition $(A,B)$. Let $C_{0}=\frac{C}{4}$ and $K=2^{3k+5}$, then there is a $K$-almost-regular induced subgraph $H_{1}$ of $G_{1}$ satisfying $e(H_{1})\geq C_{0}m^{1+1/k}$, where $m$ is the number of vertices of $H_{1}$. Also, let $H$ be the subgraph of $G$ induced by $V(H_{1})$.
    For any two distinct vertices $u,v\in V(H)$, we denote by $\mathcal{P}_{u,v}$ the set of alternating paths of length $k$ between vertices $u$ and $v$ in $H$, and let $P_{u,v}=|\mathcal{P}_{u,v}|$. Furthermore, let $A_{u,v}$ denote the number of ordered pairs of paths $(P_{1},P_{2})\in\mathcal{P}_{u,v}^{2}$ which intersect in a vertex different from $u,v$. Finally, let $B_{u,v}$ denote the number of ordered pairs of distinct paths $(P_{1},P_{2})\in\mathcal{P}_{u,v}^{2}$ such that some internal vertex of $P_{1}$ and some internal vertex of $P_{2}$ are connected by an edge of $H$.
@@
    Finally, if $H$ has no induced nondegenerate $2k$-cycles, then $\sum_{u,v}B_{u,v}\geq\frac{1}{4k}\hom(C_{2k},H_{1})$. To see why, note that every nondegenerate cycle $(x_{1},\dots,x_{2k})$ must have a chord, since it is not induced. Then, there exists a way to shift the indices of the vertices, say by setting $y_{i}=x_{i+t\bmod 2k}$ for some $t$, such that the chord goes between the vertex sets $\{y_{2},\dots,y_{k}\}$ and $\{y_{k+2},\dots,y_{2k}\}$. In this case, the cycle $(y_{1},\dots,y_{2k})$ corresponds to a pair of paths between $y_{1}$ and $y_{k+1}$ with an edge between their internal vertices, which is counted in $B_{y_{1},y_{k+1}}$. This argument shows that for any non-degenerate $2k$-cycle, one of its $2k$ cyclic relabeling corresponds to a pair of paths counted in the sum $\sum_{u,v}B_{u,v}$. Hence, $2k\sum_{u,v}B_{u,v}$ is at least the number of nondegenerate $2k$-cycles, and so $\sum_{u,v}B_{u,v}\geq\frac{1}{4k}\cdot\hom(C_{2k},H_{1})$.
@@
    Let us fix these vertices and define $\delta=2^{2k+10}\cdot{8k}/{\sqrt{C_{0}}}$. Since $B_{u,v}$ counts certain pairs of paths in $\mathcal{P}_{u,v}$, we always have $B_{u,v}\leq P_{u,v}^{2}$. Thus, we deduce that $8k\delta^{-1}A_{u,v}\leq 8kB_{u,v}\leq 8kP_{u,v}^{2}$ and so $A_{u,v}\leq\delta P_{u,v}^{2}$. Also, we directly have both $B_{u,v}\geq\frac{1}{16k}P_{u,v}^{2}$ and $8kP_{u,v}^{2}\geq 8kB_{u,v}\geq\frac{1}{2}C_{0}^{2k}$, implying $P_{u,v}\geq C_{0}^{k}/4\sqrt{k}$.
    We now define two auxiliary graphs, a red and a blue one, on the set of vertices $\mathcal{P}_{u,v}$. A pair of distinct paths $P_{1},P_{2}\in\mathcal{P}_{u,v}$ is a red edge if the paths $P_{1},P_{2}$ intersect in some of their internal vertices. Also, $P_{1}P_{2}$ is a blue edge if there exists an edge of $H$ between the internal vertices of $P_{1}$ and $P_{2}$. Let us denote the red graph by $G_{R}$ and the blue graph by $G_{B}$. Note that an ordered pair of paths $(P_{1},P_{2})$ is counted in $A_{u,v}$ precisely if $P_{1}P_{2}$ is a red edge or if $P_{1}=P_{2}$. Thus, we have that the number of red edges is precisely $e(G_{R})=(A_{u,v}-P_{u,v})/{2}\leq\delta P_{u,v}^{2}$. Similarly, we have $e(G_{B})=B_{u,v}/{2}\geq\frac{1}{32k}P_{u,v}^{2}$, since $B_{u,v}$ only counts pairs of distinct paths.
@@ 4.1 The cube graph
We finish the section by showing Proposition 1.5, which states that $\mathrm{ex}^{*}(n,Q_{8},s)\leq O_{s}(n^{8/5})$, where $Q_{8}$ denotes the graph of a three-dimensional cube. Before we start the proof, we show that almost all paths of a given length in a $K_{s,s}$-free graph are induced. Although we need this statement only for paths of length three, we prove it in full generality since we believe it might be of independent interest.
@@
    Let $G$ be an $n$ vertex graph with average degree at least $Cn^{3/5}$ containing no $K_{s,s}$. Partition the vertex set of $G$ into parts $A,B$ such that at least half of the edges cross the partition, and let $G_{1}$ be the graph of crossing edges. By Lemma 4.2 applied with $\alpha=\frac{3}{5}$, $G_{1}$ has a $K$-almost-regular induced subgraph $H_{1}$ with $m$ vertices and average degree $d$ satisfying $d\geq C_{0}m^{3/5}$, where $C_{0}=\frac{C}{4}$ and $K=2^{10}$. Let $H$ be the subgraph of $G$ induced on $V(H_{1})$.
@@
    We claim that if $C_{0}$ is sufficiently large, one has $e(U,V)\geq C_{0}(|U|+|V|)^{4/3}$, which allows us to apply Theorem 4.1 with $k=3$. To verify that there are sufficiently many edges between $U$ and $V$, we recall that $d\geq C_{0}m^{3/5}$, which implies
    $e(U,V)=\Omega(d^{3}/m)\geq\Omega\left(d^{4/3}C_{0}^{5/3}\right)\geq C_{0}\cdot(2Kd)^{4/3}\geq C_{0}(|U|+|V|)^{4/3}.$
@@ 5 Concluding remarks
Acknowledgments. We thank anonymous referees for numerous thoughtful comments and the observation that the original statement of the Conjecture 1.1 had a trivial counterexample, if $H$ is allowed to be disconnected. We note that a similar observation was made by Sheffield [50] too.